OpenStudy (anonymous):

How do I solve this express the ratio 225 : 750 in its lowest terms?

OpenStudy (jdoe0001):

just simplify \(\bf \cfrac{225}{750}\)

OpenStudy (anonymous):

Find the common factor in both.......in this case is 75.....therefore by division the simplest ratio is 3:10

OpenStudy (wolf1728):

Factor each number into its prime factors: 225 = 5*5*3*3 750 = 5*5*5*3*2 It seems we can cancel out 2 fives and 1 three from both numbers giving us 3 / 5*2 = 3/10

OpenStudy (08surya):

both the numbers are divisible by 3 so divide it